OpenStudy (anonymous):

Which equations are equivalent to the given equation? 0.4y=10 Choose all answers that are correct A. 0.4y=10 divided by 0.4 B. 0.4y divided by 0.4=10 divided by 0.4 C. 0.4y multiplied by 2.5=10 mulitiplied by 2.5 D. 0.4y divided by 0.4=10 multiplied by 0.4

OpenStudy (anonymous):

?

OpenStudy (peach3113):

B.

OpenStudy (dayakar):

A

OpenStudy (anonymous):

any others because it says choose all that aply

OpenStudy (dayakar):

0.4y/0.4 = 10/0.4 y = 10/0.4

OpenStudy (peach3113):

B and A.

OpenStudy (anonymous):

ok thanks both

OpenStudy (peach3113):

You're welcome. :P

OpenStudy (dayakar):

A is correct because u can find y value

OpenStudy (peach3113):

B is correct because you can find y value also.

OpenStudy (anonymous):

it helps because i am trying to get homework done because it is my BIRTHDAY

OpenStudy (peach3113):

*throws confetti in the air* *blows one of those loud annoying things people always have at birthdays* *hands you a gazillion balloons* HAPPY BIRTHDAY!!!!

OpenStudy (anonymous):

thank you so much! :D

OpenStudy (anonymous):

I became your friend! :) K gotta go now bye!